1. For each genotype, indicate whether it is heterozygous (HE) or homozygous (HO) AA __HO__ Bb __HE__ cc __HO__ Dd __HE__ 2. For each of the genotypes below, determine the phenotype. (P=purple, p=white) Purple flowers are dominant to white flowers PP ________________________Homogeneous Dominate Pp ________________________Heterozygous pp _______________________Homozygous Recessive 3. In guinea pigs, the allele for short hair is dominant. S= Short Hair s= long Hair What genotype would a heterozygous short haired guinea pig have? ____Ss___ What genotype would a pure-breeding short haired guinea pig have? ___SS____ What genotype would a long-haired guinea pig have? __ss______ 4. Show the cross for two heterozygous guinea pigs. What percentage of the offspring will have short hair? ___50 percent_____ What percentage of the offspring will have long hair? _____50 percent__ 6. Two short haired guinea pigs are mated several times. Out of 100 offspring, 25 of them have long hair. What are the probable genotypes of the parents?

___Ss_____ x ____Ss_______ Show the cross to prove it! 7. In peas, the yellow allele is dominant to the green allele… Y=yellow y=green 8. Imagine you breed two green pea plants together, and they have 100 offspring. 25% are green and 75% are yellow. What are the genotypes of the parent peas and the offspring peas? Yy and Yy
Answer #1

Answer:

1).

AA __HO__

Bb __HE__

cc __HO__

Dd __HE__

2. For each of the genotypes below, determine the phenotype. (P=purple, p=white) Purple flowers are dominant to white flowers

PP =Homogeneous Dominate

Pp =Heterozygous

pp =Homozygous Recessive

3).

The genotype of a heterozygous short haired guinea pig = Ss

The genotype of a pure-breeding short haired guinea pig have = SS

The genotype of a long-haired guinea pig = ss

4).

Ss x Ss –Parents

S

s

S

SS (short)

Ss (short)

S

Ss (short)

ss (long)

Short hair = SS + Ss = ¾ = 75%

Long hair = ss = 25%

6). Ss x Ss—Parents

Ss x Ss –Parents

S

s

S

SS (short)

Ss (short)

S

Ss (short)

ss (long)

Short hair = SS + Ss = ¾ = 75%

Long hair = ss = 25%

7. In peas, the yellow allele is dominant to the green allele… Y=yellow y=green

8. Imagine you breed two green pea plants together, and they have 100 offspring.

yy x yy –Parents

yy –Progeny

All progeny are green and no yellow progeny.

-----------------

25% are green and 75% are yellow. What are the genotypes of the parent peas and the offspring peas= Yy and Yy

Ss x Ss –Parents

Y

Y

Y

YY (yellow)

Yy (ywellow)

y

Yy (yellow)

yy (green)

Yellow = YY + Yy = ¾ = 75%

Green = yy = 25%